OpenStudy (anonymous):

What is the missing integer in this problem? (Do not show work or skip spaces when answering this question. Make sure to type in only the missing integer) 27 - ____ = -9

OpenStudy (anonymous):

27-x=-9 add 27 to both sides

OpenStudy (ajprincess):

Let the missing integer be x.\. Then 27-x=-9 x=27+9 Can u do nw?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

36

OpenStudy (ajprincess):

ya.:)